# Self-Dual Codes and LCD Codes in Sum-Rank Metric

### v1

@@ 2.3 Sum-Rank Codes
In this subsection, we recall some knowledge about sum-rank codes which play crucial roles in our work. We begin with the definition of sum-rank metric as well as those related notions from [7]. Let $\mathbb{F}_{q}^{(m,n)}$ be the set of all $m\times n$ matrices over $\mathbb{F}_{q}$. This is a linear space over $\mathbb{F}_{q}$ of dimension $mn$. Let $t,\,m_{1},\cdots,m_{t},\,n_{1},\cdots,\,n_{t}$ be positive integers such that $m_{i}\leq n_{i}$ for all $i\in[t]$ and $n_{1}\geq\cdots\geq n_{t}$ where $[t]=\{1,\ldots,t\}$. We consider the product of $t$ matrix spaces
@@
A subset $C$ of the linear space $\mathbb{F}_{q}^{(m_{1},n_{1}),\ldots,(m_{t},n_{t})}$ with the sum-rank metric is called a sum-rank code with block length $t$ and matrix sizes $m_{1}\times n_{1},\ldots,m_{t}\times n_{t}$. Its minimum sum-rank distance is defined as
$d_{sr}(C)=\min\limits_{M\neq N,\,M,N\in C}d_{sr}(M,N).$
Its minimum sum-rank weight is defined as
$wt_{sr}(C)=\min\limits_{M\neq\bm{0},M\in C}wt_{sr}(M).$
@@
If $C$ is a linear sum-rank code, $d_{sr}(C)=wt_{sr}(C)$.
@@
Let $C\subseteq\mathbb{F}_{q^{m}}^{N}$. Then the minimum sum-rank distance of $C$ is $d_{sr}(C)=\min\limits_{\bm{X}\neq\bm{Y},\,\bm{X},\bm{Y}\in C}d_{sr}(\bm{X},\bm{Y})$. It is easy to prove that the definition is the same as that above when it comes to the special class of sum-rank codes. For more details about the class of sum-rank codes, readers can refer to [48], [49], [51], [52] and so on.
@@
A sequence of sum-rank codes $C_{1}$, $C_{2}$, $\cdots$ with matrix size $m\times n$ is said to be asymptotically good if the block length $t_{i}$ of $C_{i}$ goes to infinity and, for $i=1,2,\ldots$, both the rate $R_{sr}(C_{i})$ and relative minimum sum-rank distance $\Delta_{sr}(C_{i})$ are positively bounded from below.
According to [15], a sum-rank code $C$ with block length $t$ and matrix size $m\times n$ is called a cyclic sum-rank code, if $M=(M_{1},\ldots,M_{t})$ is a codeword in $C$, $M_{i}\in\mathbb{F}_{q}^{(m,n)}$, $i=1,\ldots,t$, implies that $(M_{t},M_{1},\ldots,M_{t-1})$ is also a codeword in $C$.
Let $M=(M_{1},\ldots,M_{t}),\,N=(N_{1},\ldots,N_{t})\in\mathbb{F}_{q}^{(m_{1},n_{1}),\ldots,(m_{t},n_{t})}$. The trace inner product of them is defined as
@@
Now we review an important construction of sum-rank codes from [12]. The author noted that the linear space $\mathbb{F}_{q}^{(m,m)}$ of all $m\times m$ matrices over $\mathbb{F}_{q}$ can be identified with the space of all $q-polynomials$,
@@
Let $C_{i}$ be a linear code in $\mathbb{F}_{q^{m}}^{t}$, $10\leq i\leq m-$ 1. Then the linear sum-rank code $SR(C_{0},\ldots,C_{m-1})$ over $\mathbb{F}_{q}$ with block length $t$ and matrix size $m\times m$.
[…]
@@
When $C_{i}\subseteq\mathbb{F}_{q^{m}}^{t}$ is a cyclic code, $10\leq i\leq m-$ 1, it is clear that $SR(C_{0},\ldots,C_{m-1})$ is a cyclic sum-rank code.
@@
([15, Theorem 3.1 and Proposition 3.1]) Let $C_{i}\subseteq\mathbb{F}_{q^{m}}^{t}$ be a $[t,k_{i},d_{i}]_{q^{m}}$-linear code, $10\leq i\leq m-$ 1. Then $SR(C_{0},\ldots,C_{m-1})$ is a linear sum-rank code over $\mathbb{F}_{q}$ with block length $t$, matrix size $m\times m$ and dimension $m(k_{0}+\cdots+k_{m-1})$. The minimum sum-rank distance $d_{sr}$ of $SR(C_{0},\ldots,C_{m-1})$ satisfies
@@
If $C\subseteq\mathbb{F}_{q}^{(m_{1},n_{1}),\ldots,(m_{t},n_{t})}$ is a sum-rank code with $M$ codewords and the minimum sum-rank distance $d_{sr}$ and there is no sum-rank code in $\mathbb{F}_{q}^{(m_{1},n_{1}),\ldots,(m_{t},n_{t})}$ with $M$ codewords and the minimum sum-rank distance $d_{sr}+1$, we call $C$ a distance-optimal sum-rank code. The authors of [15] constructed a family of distance-optimal cyclic sum-rank codes in $\mathbb{F}_{2}^{(2,2),\ldots,(2,2)}$ using the method above.
Lemma 2.16.
([15, Corollary 7.3]) Let $l\geq 3$, $C_{0}$ be the cyclic code of length $14^{l}-$ 1 over $\mathbb{F}_{4}$ with generator polynomial $x-1$ and $C_{1}$ be the BCH code $C_{(4,4^{l}-1,4,0)}$. Then $SR(C_{0},C_{1})\subseteq\mathbb{F}_{2}^{(2,2),\ldots,(2,2)}$ has dimension $4(4^{l}-l-2)$ and minimum sum-rank distance $4$. Furthermore, the sum-rank cyclic code $SR(C_{0},C_{1})$ is distance-optimal.
3 Characterization of Dual Codes of Some Sum-Rank Codes
In this section, we study two classes of sum-rank codes and characterize their dual codes.
3.1 The Dual Code of $SR(C_{0},C_{1})$
In this subsection, the dual code of $SR(C_{0},C_{1})$ is characterized. In order to make it more clear, we review the method of constructing sum-rank codes in [12] mentioned above once again, but in this subsection, we only consider the particular case where both $q$ and $m$ are equal to $2$.
Firstly, we have the isomorphism
$\mathbb{F}_{2}^{(2,2)}\cong\{a_{0}x+a_{1}x^{2}\,:\,a_{0},\,a_{1}\in\mathbb{F}_{4}\}.$
Let $C_{0}$, $C_{1}$ be linear codes over $\mathbb{F}_{4}$ of length $t$,
$SR(C_{0},C_{1})=\{SR(\bm{c}_{0},\bm{c}_{1})\,:\,\bm{c}_{0}=(c_{0,1},\ldots,c_{0,t})\in C_{0},\bm{c}_{1}=(c_{1,1},\ldots,c_{1,t})\in C_{1}\},$
where
$SR(\bm{c}_{0},\bm{c}_{1})=(c_{0,1}x+c_{1,1}x^{2},\ldots,c_{0,t}x+c_{1,t}x^{2}).$
$SR(C_{0},C_{1})$ is a linear sum-rank code over $\mathbb{F}_{2}$ with block length $t$ and matrix size $2\times 2$.
We have the key lemma below.
Lemma 3.1.
@@ 3.1 The Dual Code of $SR(C_{0},C_{1})$
Let $C_{i}$ be a linear code in $\mathbb{F}_{q^{m}}^{t}$, $10\leq i\leq m-$ 1. It is very challenging to prove the more general result:
@@ 3.2 The Dual Code of $Mat_{\mathcal{B}}(C)$
In this subsection, we aim to characterize the dual code of another class of sum-rank codes. It is inspired from [64] which mainly studied the duality theory in rank metric. In Remark 2.10, we point out that $\mathbb{F}_{q}^{(m_{1},n_{1}),\ldots,(m_{t},n_{t})}\cong\mathbb{F}_{q^{m}}^{N}$ if $m_{1}=\cdots=m_{t}=m$ and $N=n_{1}+\cdots+n_{t}$. This is an $\mathbb{F}_{q}$-linear isomorphism between the two linear spaces. Now, we give the isomorphism exactly. Let =$\mathcal{B}=\{\gamma_{1},\ldots,\gamma_{m}\}$ be the basis of $\mathbb{F}_{q^{m}}$ over $\mathbb{F}_{q}$ and $\bm{X}=(\bm{X}^{(1)},\ldots,\bm{X}^{(t)})\in\mathbb{F}_{q^{m}}^{N}$, where $\bm{X}^{(i)}=(\bm{X}^{(i)}_{1},\ldots,\bm{X}^{(i)}_{n_{i}})\in\mathbb{F}_{q^{m}}^{n_{i}}$. Set $Mat_{\mathcal{B}}^{(i)}\,:\,\mathbb{F}_{q^{m}}^{n_{i}}\rightarrow\mathbb{F}_{q}^{(m,n_{i})}$, $Mat_{\mathcal{B}}^{(i)}(\bm{X}^{(i)})=M_{\mathcal{B}}^{(i)}$, where $\bm{X}_{s}^{(i)}=\sum_{j=1}^{m}(M_{\mathcal{B}}^{(i)})_{js}\gamma_{j}$, $1\leq s\leq n_{i}$. Let $Mat_{\mathcal{B}}\,:\,\mathbb{F}_{q^{m}}^{N}\rightarrow\mathbb{F}_{q}^{(m,n_{1}),\ldots,(m,n_{t})}$, $Mat_{\mathcal{B}}(\bm{X})=(Mat_{\mathcal{B}}^{(1)}(\bm{X}^{(1)}),\ldots,Mat_{\mathcal{B}}^{(t)}(\bm{X}^{(t)}))=(M_{\mathcal{B}}^{(1)},\ldots,M_{\mathcal{B}}^{(t)})=M_{\mathcal{B}}$. We note that these notations will be used throughout the subsection.
Remark 3.3.
Note that the $s$-th row of $M_{\mathcal{B}}^{(i)}$ is the expansion of $X_{s}^{(i)}$ over the basis $\mathcal{B}$.
The following result is immediate.
@@
Let Trace : $\mathbb{F}_{q^{m}}\,\rightarrow\,\mathbb{F}_{q}$ be the $\mathbb{F}_{q}$-linear trace map given by $Trace(\alpha)\,:=\alpha+\alpha^{q}+\cdots+\alpha^{q^{m-1}}$ for all $\alpha\in\mathbb{F}_{q^{m}}$. The bases =$\mathcal{B}=\{\gamma_{1},\ldots,\gamma_{m}\}$ and =$\mathcal{B}^{\prime}=\{\gamma_{1}^{\prime},\ldots,\gamma_{m}^{\prime}\}$ of $\mathbb{F}_{q^{m}}$ over $\mathbb{F}_{q}$ are said to be dual if $Trace(\gamma_{i}\gamma_{j}^{\prime})=\delta_{ij}$ for all $i,\,j\in\{1,\ldots,m\}$. From [42, p58], we get the fact that for every basis $\mathcal{B}$ of $\mathbb{F}_{q^{m}}$ over $\mathbb{F}_{q}$ there exists a unique dual basis $\mathcal{B}^{\prime}$. In the following, we give another key lemma with the definition.
Lemma 3.6.
Let $C$ be a linear code in $\mathbb{F}_{q^{m}}^{N}$ and =$\mathcal{B}=\{\gamma_{1},\ldots,\gamma_{m}\}$ and =$\mathcal{B}^{\prime}=\{\gamma_{1}^{\prime},\ldots,\gamma_{m}^{\prime}\}$ be dual bases of $\mathbb{F}_{q^{m}}$ over $\mathbb{F}_{q}$. We have
@@
Thus, $M_{\mathcal{B}^{\prime}}\in Mat_{\mathcal{B}}(C)^{\bot_{tr}}$, giving $Mat_{\mathcal{B}^{\prime}}(C^{\bot})\subseteq Mat_{\mathcal{B}}(C)^{\bot_{tr}}$.
Let $k=\dim_{\mathbb{F}_{q^{m}}}(C)$. Then
@@
Let the bases =$\mathcal{B}=\{\gamma_{1},\ldots,\gamma_{m}\}$ and =$\mathcal{B}^{\prime}=\{\gamma_{1}^{\prime},\ldots,\gamma_{m}^{\prime}\}$ of $\mathbb{F}_{q^{m}}$ be dual bases of $\mathbb{F}_{q^{m}}$ over $\mathbb{F}_{q}$. If $\mathcal{B}=\mathcal{B}^{\prime}$, $\mathcal{B}$ is said to be a self-dual basis. By [35, Theorem 1], a self-dual basis $\mathcal{B}$ of $\mathbb{F}_{q^{m}}$ over $\mathbb{F}_{q}$ exists if and only if $q$ is even or both $q$ and $m$ are odd. Using these knowledge, it is easy to get the corollary below.
@@ 4 Self-Dual Codes and LCD Codes in Sum-Rank Metric
In this section, we firstly give the definition of self-dual sum-rank codes and LCD sum-rank codes and derive some properties of such codes.
@@
In the following, we give two methods of constructing self-dual sum-rank codes and LCD sum-rank codes from Euclidean self-dual codes and Euclidean LCD codes.

### v2

@@ 2.3 Sum-Rank Codes
In this subsection, we review some knowledge about sum-rank codes which play crucial roles in our work. We begin with the definition of the sum-rank metric as well as those related notions from [7]. Let $\mathbb{F}_{q}^{(m,n)}$ be the set of all $m\times n$ matrices over $\mathbb{F}_{q}$. This is a linear space over $\mathbb{F}_{q}$ of dimension $mn$. Let $t,\,m_{1},\cdots,m_{t},\,n_{1},\cdots,\,n_{t}$ be positive integers such that $m_{i}\leq n_{i}$ for all $i\in[t]$ and $n_{1}\geq\cdots\geq n_{t}$ where $t]=\{1,\ldots,t\}$. We consider the product of $t$ matrix spaces
@@
A subset $C$ of the linear space $\mathbb{F}_{q}^{(m_{1},n_{1}),\ldots,(m_{t},n_{t})}$ with the sum-rank metric is called a sum-rank code with block length $t$ and matrix sizes $m_{1}\times n_{1},\ldots,m_{t}\times n_{t}$. Its minimum sum-rank distance and minimum sum-rank weight are defined as
$d_{sr}(C)=\min\limits_{M\neq N,\,M,N\in C}d_{sr}(M,N)\ \text{and}\ wt_{sr}(C)=\min\limits_{M\neq\bm{0},M\in C}wt_{sr}(M).$
@@
Obviously, if $C$ is a linear sum-rank code, $d_{sr}(C)=wt_{sr}(C)$.
@@
Let $C\subseteq\mathbb{F}_{q^{m}}^{N}$. Then the minimum sum-rank distance of $C$ is $d_{sr}(C)=\min\limits_{\bm{X}\neq\bm{Y},\,\bm{X},\bm{Y}\in C}d_{sr}(\bm{X},\bm{Y})$. It is easy to prove that the definition is the same as that above when it comes to the special class of sum-rank codes. For more details about the class of sum-rank codes, readers can refer to [36], [38], [39] and so on.
@@
A sequence of sum-rank codes $C_{1}$, $C_{2}$, $\cdots$ with matrix size $m\times n$ is said to be asymptotically good if for $i=1,2,\ldots$, the block length $t_{i}$ of $C_{i}$ goes to infinity and both the rate $R_{sr}(C_{i})$ and the relative minimum sum-rank distance $\Delta_{sr}(C_{i})$ are positively bounded from below.
In addition, according to [12], a sum-rank code $C$ with block length $t$ and matrix size $m\times n$ is called a cyclic sum-rank code, if $M=(M_{1},\ldots,M_{t})$ is a codeword in $C$ where $M_{i}\in\mathbb{F}_{q}^{(m,n)}$, $i=1,\ldots,t$ implies that $(M_{t},M_{1},\ldots,M_{t-1})$ is also a codeword in $C$.
We end up this subsection with the definition of the dual code of a sum-rank code which can be found in [7]. Let $M=(M_{1},\ldots,M_{t}),\,N=(N_{1},\ldots,N_{t})\in\mathbb{F}_{q}^{(m_{1},n_{1}),\ldots,(m_{t},n_{t})}$. The trace inner product of them is defined as
@@
This definition is the starting point of our work.
3 Characterization of Dual Codes of Some Sum-Rank Codes
In this section, we research two classes of sum-rank codes and characterize their dual codes.
3.1 The Dual Code of $SR(C_{0},C_{1})$
Firstly, we review an essential construction of sum-rank codes from [10]. The author noted that the linear space $\mathbb{F}_{q}^{(m,m)}$ of all $m\times m$ matrices over $\mathbb{F}_{q}$ can be identified with the space of all qpolynomialsq-polynomials,
@@ 3.1 The Dual Code of $SR(C_{0},C_{1})$
Let $C_{i}$ be a linear code in $\mathbb{F}_{q^{m}}^{t}$, $0\leq i\leq m-1$. Define the linear sum-rank code $SR(C_{0},\ldots,C_{m-1})$ over $\mathbb{F}_{q}$ with block length $t$ and matrix size $m\times m$ as
$SR(C_{0},\ldots,C_{m-1})=\{SR(\bm{c}_{0},\ldots,\bm{c}_{m-1})\,:\,\bm{c}_{i}=(c_{i,1},\ldots,c_{i,t})\in C_{i},\,0\leq i\leq m-1\},$
@@
When $C_{i}\subseteq\mathbb{F}_{q^{m}}^{t}$ is a cyclic code, $0\leq i\leq m-1$, it is clear that $SR(C_{0},\ldots,C_{m-1})$ is a cyclic sum-rank code.
@@
([12, Theorem 3.1 and Proposition 3.1]) Let $C_{i}\subseteq\mathbb{F}_{q^{m}}^{t}$ be a $t,k_{i},d_{i}]_{q^{m}}-$linear code, $0\leq i\leq m-1$. Then $SR(C_{0},\ldots,C_{m-1})$ is a linear sum-rank code over $\mathbb{F}_{q}$ with block length $t$, matrix size $m\times m$ and dimension $m(k_{0}+\cdots+k_{m-1})$. The minimum sum-rank distance $d_{sr}$ of $SR(C_{0},\ldots,C_{m-1})$ satisfies
@@
In this subsection, we only consider the particular case where both $q$ and $m$ are equal to $2$. In the following, we characterize the dual code of $SR(C_{0},C_{1})$.
Theorem 3.4.
@@
Let $C_{i}$ be a linear code in $\mathbb{F}_{q^{m}}^{t}$, $0\leq i\leq m-1$. It is very challenging to prove the more general result:
@@ 3.2 The Dual Code of $Mat_{\mathcal{B}}(C)$
In this subsection, we aim to characterize the dual code of another class of sum-rank codes. It is inspired from [48] which mainly studied the duality theory in rank metric. In Remark 2.7, we point out that $\mathbb{F}_{q}^{(m_{1},n_{1}),\ldots,(m_{t},n_{t})}\cong\mathbb{F}_{q^{m}}^{N}$ if $m_{1}=\cdots=m_{t}=m$ and $N=n_{1}+\cdots+n_{t}$. This is an $\mathbb{F}_{q}$-linear isomorphism between the two linear spaces. Now, we give the isomorphism exactly. Let $\mathcal{B}=\{\gamma_{1},\ldots,\gamma_{m}\}$ be the basis of $\mathbb{F}_{q^{m}}$ over $\mathbb{F}_{q}$ and $\bm{X}=(\bm{X}^{(1)},\ldots,\bm{X}^{(t)})\in\mathbb{F}_{q^{m}}^{N}$, where $\bm{X}^{(i)}=(\bm{X}^{(i)}_{1},\ldots,\bm{X}^{(i)}_{n_{i}})\in\mathbb{F}_{q^{m}}^{n_{i}}$. Set $Mat_{\mathcal{B}}^{(i)}\,:\,\mathbb{F}_{q^{m}}^{n_{i}}\rightarrow\mathbb{F}_{q}^{(m,n_{i})}$, $Mat_{\mathcal{B}}^{(i)}(\bm{X}^{(i)})=M_{\mathcal{B}}^{(i)}$, where $\bm{X}_{s}^{(i)}=\sum_{j=1}^{m}(M_{\mathcal{B}}^{(i)})_{js}\gamma_{j}$, $1\leq s\leq n_{i}$. In other words, the $s$-th row of $M_{\mathcal{B}}^{(i)}$ is the expansion of $\bm{X}_{s}^{(i)}$ over the basis $\mathcal{B}$. Let $Mat_{\mathcal{B}}\,:\,\mathbb{F}_{q^{m}}^{N}\rightarrow\mathbb{F}_{q}^{(m,n_{1}),\ldots,(m,n_{t})}$, $Mat_{\mathcal{B}}(\bm{X})=(Mat_{\mathcal{B}}^{(1)}(\bm{X}^{(1)}),\ldots,Mat_{\mathcal{B}}^{(t)}(\bm{X}^{(t)}))=(M_{\mathcal{B}}^{(1)},\ldots,M_{\mathcal{B}}^{(t)})=M_{\mathcal{B}}$. We note that these notations will be used throughout the subsection. The following result is immediate.
@@
Let Trace : $\mathbb{F}_{q^{m}}\,\rightarrow\,\mathbb{F}_{q}$ be the $\mathbb{F}_{q}$-linear trace map given by $Trace(\alpha)\,:=\alpha+\alpha^{q}+\cdots+\alpha^{q^{m-1}}$ for all $\alpha\in\mathbb{F}_{q^{m}}$. The bases $\mathcal{B}=\{\gamma_{1},\ldots,\gamma_{m}\}$ and $\mathcal{B}^{\prime}=\{\gamma_{1}^{\prime},\ldots,\gamma_{m}^{\prime}\}$ of $\mathbb{F}_{q^{m}}$ over $\mathbb{F}_{q}$ are said to be dual if $Trace(\gamma_{i}\gamma_{j}^{\prime})=\delta_{ij}=\begin{cases}1,&\mbox{if }i=j,\\ 0,&\mbox{if }i\neq j,\end{cases}$ for all $i,\,j\in\{1,\ldots,m\}$. From [30, p58], we get the fact that for every basis $\mathcal{B}$ of $\mathbb{F}_{q^{m}}$ over $\mathbb{F}_{q}$ there exists a unique dual basis $\mathcal{B}^{\prime}$. In the following, we give the characterization of $Mat_{\mathcal{B}}(C)^{\bot_{tr}}$.
Theorem 3.8.
Let $C$ be a linear code in $\mathbb{F}_{q^{m}}^{N}$ and $\mathcal{B}=\{\gamma_{1},\ldots,\gamma_{m}\}$ and $\mathcal{B}^{\prime}=\{\gamma_{1}^{\prime},\ldots,\gamma_{m}^{\prime}\}$ be dual bases of $\mathbb{F}_{q^{m}}$ over $\mathbb{F}_{q}$. We have
@@
Thus, $M_{\mathcal{B}^{\prime}}\in Mat_{\mathcal{B}}(C)^{\bot_{tr}}$, giving $Mat_{\mathcal{B}^{\prime}}(C^{\bot})\subseteq Mat_{\mathcal{B}}(C)^{\bot_{tr}}$. Let $k=\dim_{\mathbb{F}_{q^{m}}}(C)$. Then
@@
Let $\mathcal{B}=\{\gamma_{1},\ldots,\gamma_{m}\}$ and $\mathcal{B}^{\prime}=\{\gamma_{1}^{\prime},\ldots,\gamma_{m}^{\prime}\}$ be dual bases of $\mathbb{F}_{q^{m}}$ over $\mathbb{F}_{q}$. If $\mathcal{B}=\mathcal{B}^{\prime}$, $\mathcal{B}$ is said to be a self-dual basis. By [23, Theorem 1], a self-dual basis $\mathcal{B}$ of $\mathbb{F}_{q^{m}}$ over $\mathbb{F}_{q}$ exists if and only if $q$ is even or both $q$ and $m$ are odd. Using these knowledge, it is easy to get the corollary below.
@@ 4 Self-Dual Codes and LCD Codes in Sum-Rank Metric
In this section, we firstly give the definitions of self-dual sum-rank codes and LCD sum-rank codes and derive some properties of such codes.
@@
Next, we give two methods of constructing self-dual codes and LCD codes in sum-rank metric from Euclidean self-dual codes and Euclidean LCD codes.